OpenStudy (suwhitney):

Find the vertical and horizontal asymptote(s), if any, of the following function...

OpenStudy (suwhitney):

\[f(x)=\frac{ (6x^2+x+12) }{(3x^2−5x+2)(x^2+2) }\]

OpenStudy (suwhitney):

So very confused :/ Please help..

OpenStudy (zugzwang):

It would have a vertical asymptote at the line x = a if \[\lim_{x \rightarrow a}f(x)=\pm \infty\]

OpenStudy (zugzwang):

Possibly for values of x where the denominator (but not the numerator) is zero.

OpenStudy (suwhitney):

Do you mind kinda working it out with me because I have no clue how to get the answer..

OpenStudy (zugzwang):

Sure :) The denominator is (3x^2 - 5x + 2)(x^2 + 2) Now when is this equal to zero?

OpenStudy (suwhitney):

Thanks, and Im not sure do I have to set both of them equal to zero?

OpenStudy (zugzwang):

No... ab = 0 means a = 0 or b = 0, only one of then need be zero besides... (x^2 + 2) is never zero for any real values of x. So I suggest you look at the other factor :)

OpenStudy (suwhitney):

So..? \[(3x^2 - 5x + 2)=0\]

OpenStudy (zugzwang):

Yes..., so can you work it out?

OpenStudy (suwhitney):

Yes.. just a second

OpenStudy (suwhitney):

Ok, so is it x=1 and x=2/3?

OpenStudy (zugzwang):

ok, good. Now, we know the denominator goes to zero as x goes to 1 and as x goes to 2/3, where does the numerator go? Does it go to zero as well? (Evaluate the numerator at x = 1 and at x = 2/3)

OpenStudy (suwhitney):

So instead of making the numerator = to 0. I replace the x's with 1, and then 2/3 and see if that = to zero?

OpenStudy (zugzwang):

Yes. If they happen to be equal to 0 at 1 or at 2/3, then you may not have an asymptote at either. If they are nonzero, then you have a vertical asymptote at x = 1 and x = 2/3

OpenStudy (suwhitney):

No neither one will make it = to zero.. is that right?

OpenStudy (zugzwang):

Yes.

OpenStudy (suwhitney):

Ok :) What's next?

OpenStudy (suwhitney):

Oh or was that it?

OpenStudy (zugzwang):

Well, it means you have vertical asymptotes at x=1 and x=2/3 Now find the horizontal asymptotes

OpenStudy (suwhitney):

Well since there isnt a horizontal is it just y=0?

OpenStudy (zugzwang):

What do you mean by "there isn't a horizontal" ?

OpenStudy (suwhitney):

The horizontal asymptote.. since the numerator couldn't = to zero?

OpenStudy (zugzwang):

If the numerator is zero, then you've simply found a root of the function (A root is a point where your function intersects the x axis)

OpenStudy (zugzwang):

Equating the numerator to 0 is not going to get you your horizontal asymptote :)

OpenStudy (suwhitney):

Oh okay :) So how do I go about finding the horizontal asymptote them?

OpenStudy (zugzwang):